OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 50 miles per hour (mph), and the other car is traveling 40 mph. How long will it take before the cars are 360 miles apart? 4 hrs 6 hrs 2 hrs 5 hrs

OpenStudy (anonymous):

D1=S1 x T and D2=S2 x T Now plug in these two expressions: S1 x T + S2 x T = 360 T(S1+S2) = 360 T=360/(S1+S2)=360/90= ITS A